OpenStudy (anonymous):

Find the direction cosines of the vector -2i + j -5k.

OpenStudy (anonymous):

Ok, first find the magnetude of this verctor: v = -2i + j - 5k |v| = sqrt((-2)^2 + 1^2 + (-5)^2) |v| = sqrt(4 + 1 + 25) |v| = sqrt(30) now, let's find the cosines, just make inner products: v . i = |v||i|cos(o1) cos(o1) = -2/sqrt(30) v . j = |v||j|cos(o2) cos(o2) = 1/sqrt(30) v . k = |v||k|cos(o3) cos(o3) = -5/sqrt(30)

OpenStudy (anonymous):

Okay I got you till finding the magnitude of the vector, what does o1, o2, o3 mean? is it possible for you to create an equation for me?

OpenStudy (anonymous):

o1, o2 and o3 are the angles that this given vector makes with the x, y and z axis respectively

OpenStudy (anonymous):

So -2 being the cosine of i we multiply it with the |v|? that's it?

OpenStudy (anonymous):

No, we must use the definition of inner product, given a vector v = ai + bj + ck, then the directors angles are defined as: \[\vec{v} \cdot î =|{\vec{v}}||î|\cos(\theta_x)\] \[\vec{v} \cdot j =|{\vec{v}}||j|\cos(\theta_y)\] \[\vec{v} \cdot k =|{\vec{v}}||k|\cos(\theta_z)\] \[|\vec{v}| = \sqrt{a^2 + b^2 + c^2}\]

OpenStudy (anonymous):

for example: \[\vec{v} \cdot i = (ai + bj + ck)\cdot 1i + 0j + 0k = 1\times a + b\times 0+ c\times 0 = a\]

OpenStudy (anonymous):

that is a generalized way of solving this

OpenStudy (anonymous):

just apply this theory to your particular case, as i did above, and you find this cosines

OpenStudy (anonymous):

Alright hang on, I'll try some more examples, it'll help me understand. I'll write down what you said, gimme 5min

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

So if I have p = i + j + k The direction cosine of p will be |p| = sqrt/1^2 + 1^2 + 1^2 It comes to the form of 1/sqrt/3

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

Thank you @M4thM1nd your really sweet :)

OpenStudy (anonymous):

:)